# Failures to Surface Harmful Contents in Video Large Language Models

### v1

@@ VI-C Detailed Prompt Analysis on Failure Cases
VII Discussion
Potential Mitigations. Several directions may help mitigate the omission of harmful content in VideoLLMs caused by design flaws. One promising direction is to improve the frame sampling strategy. Recent studies explore new approaches such as selecting frames based on visual-textual relevance to the input prompt [37], which may, to some extent, increase the likelihood of capturing harmful segments that would otherwise be omitted. Another mitigation strategy is to incorporate auxiliary image-level checks before video understanding. Specifically, randomly selected frames can be passed through a pretrained MLLM to identify harmful content. Since these models typically do not employ token compression, they are better at detecting fine-grained harmful signals. However, the added computational burden may significantly reduce its practicality. Finally, increasing the weight of visual representations during modality fusion may reduce the dominance of textual features in the final output, thereby improving the model’s responsiveness to visual cues, including harmful content.
Long Video Understanding. Recent advances have introduced VideoLLMs designed for long video understanding (tens of minutes to hours) [38, 39]. However, the key design flaws identified in this paper still persist. For example, these models continue to rely on sparse temporal sampling, which leaves them vulnerable to the same omission issues observed in shorter videos. As shown in our analysis of harmful clip duration in Appendix, the minimum duration required for a harmful segment to evade all sampled frames increases with video length. This scaling effect makes it even easier to insert undetected harmful content in long-form videos, thereby posing greater security risks. Given the high deployment cost and the current immaturity of long VideoLLMs, we leave a detailed investigation of their safety properties to future work.
Proprietary Models. Although our study focuses on open-source VideoLLMs, the revealed design flaws may still persist in proprietary MLLMs [40] due to similar preprocessing and architectures. For example, Gemini 1.5-Pro also adopts uniform sampling of 16 frames per video [41]. A thorough investigation of harmful content omission in proprietary models is left for future work.
VIII Conclusion
This work identifies and systematically analyzes three fundamental design flaws in current VideoLLMs: sparse uniform sampling, which leaves large portions of the video unchecked; token under-sampling, which leads to the loss of localized spatial information; and modality fusion imbalance, which suppresses visual signals even when harmful content is captured by the encoder. To demonstrate the consequences of these flaws, we propose three zero-query, black-box attacks that insert harmful content through frame replacement, picture-in-picture, and transparent overlays. Despite the content being readily noticeable to human viewers, these attacks consistently achieve high Harmful Omission Rates across multiple mainstream VideoLLMs. This work serves as an early step toward understanding the structural vulnerabilities of VideoLLMs in open-world, safety-critical scenarios. We call for rethinking core design choices and building models that are not only accurate but also safe and reliable.
Appendix
A Overview of Mainstream VideoLLMs
Table III summarizes the strategies of frame selection and token compression in mainstream VideoLLMs. As aligned with our core analysis in the main text, most existing models employ uniform frame sampling, either by selecting a fixed number of frames (e.g., 16 or 32) or by applying a low and fixed sampling rate. This design significantly facilitates adversarial attempts to insert harmful content, as large portions of the video are systematically ignored. While a few models incorporate additional frame selection mechanisms, such as keyframe selection [11] or KNN-based clustering [28], the total number of selected frames remains small. Thus, our core conclusion regarding the vulnerability of sparse sampling remains unaffected.
TABLE III: Overview of frame selection and token compression strategies in mainstream VideoLLMs. FPS: frames per second.
Year Model Frame Selection Token Compression
2024 LLaMA-VID [27] 1 FPS pooling, 2 tokens per frame
2024 LLaVA-Hound [42] 10 not mentioned
2024 VideoLLaVA [43] 8 not mentioned
2024 VideoLLaMA 2 [10] 8/16 3D Conv & 3D Pool
2024 LLaVA-OneVision [6] 32 bilinear interpolation
2024 LLaVA-Video [8] 1 FPS average pooling
2024 LLaVA-NeXT-Video [9] 4/8/16/32/64 bilinear interpolation
2024 ShareGPT4Video [11] 16 not mentioned
2024 Chat-UniVi [28] KNN average pooling
2025 Apollo [44] 2 FPS 16 tokens per frame
2025 NVILA [22] 256 $\rightarrow$ 8 pooling
2025 ViLaMP [23] 1 FPS differential feature merging
2025 VideoLLaMA 3 [26] 1 FPS bilinear interpolation
2024 LongVLM [3] 100 hierarchical token merging
2024 VideoStreaming [45] 16 average pooling
2025 AKS [25] 1 FPS not mentioned
2025 FRAG [46] 256 $\rightarrow$ 32 not mentioned
B Hyperparameter Analyses
As shown in Figure 7, the probability of missing the inserted segment is high when either the sampling frame count is low or the replaced segment is short. Our results show that when only 4 or 8 frames are sampled, the inserted content is always captured by at most one frame, regardless of duration (from 1% to 10%). In fact, for the inserted segment to appear in more than one sampled frame, its duration must exceed the sampling interval. This becomes particularly critical for VideoLLMs that sample 16 frames, a configuration commonly adopted and often regarded as optimal for minute-long videos [23]. Under this setting, any segment shorter than 6% of the video duration is guaranteed to be sampled by at most one frame. For a one-minute video, a 4-second harmful segment is sufficient to meet that requirement. Also, such a segment is perceptually salient to human viewers, but is often omitted by VideoLLMs. The analysis above also justifies our experimental setting of $t_{r}=4$ in the main evaluation.
Figure 8: Attack performance under different PiP scaling ratios.

### v2

@@ Acknowledgments
Table 3: Examples of failure cases under more detailed prompt. The model correctly detects the presence of harmfulness but fails to localize or describe it accurately. R: Real information. M: Model output. Red denotes that the model output is inaccurate.
@@
Detailed Prompt Analysis on Failure Cases
In our experiments, we use HOR as the primary metric to assess whether the model could detect harmful content. However, further analysis reveals that even in cases where attacks are deemed unsuccessful (i.e., the VideoLLM provides a positive response to our prompt), the model still struggles when presented with more detailed question regarding the harmful content. In particular, we query the model with the following prompt: “Please check if the video contains crime. If yes, specify the time range (start and end) in minutes and seconds, spatial location (one of center, corner, full frame), and briefly describe the scene.”
Table 3 presents several representative examples. We find that the model often fails to provide accurate responses regarding the temporal occurrence, spatial location, or the nature of the harmful content. Specifically, in all cases, the VideoLLM is unable to correctly identify the time range where the harmful content occurs. More surprisingly, many time ranges output by the model exceed the actual duration of the video, indicating that the model cannot temporally locate specific segments. This limitation is understandable, as the model operates on sparsely sampled frames and therefore lacks coherent access to the temporal sequence. For spatial localization, the model performs relatively well when identifying harmful content located in the corner. However, it struggles to distinguish between full-frame and center regions. This limitation can be attributed to excessive token compression applied during visual encoding, which tends to blur fine-grained spatial information and reduce the model’s ability to preserve positional distinctions. As a result, the model’s spatial localization ability remains coarse and lacks precision. Regarding scene-level descriptions of harmful content, the model frequently produces responses that do not match the inserted harmful content at all. In most cases, the description corresponds to benign elements from the original video (which, notably, the model does not consider harmful when provided alone), or it refers to hallucinatory content that never appears in the video, such as in Case 3 of PPA.